If (x - 4) is a factor of x² - x - w = 0, then the value of w is:

Answer by stanbon(75887)

If (x - 4) is a factor of x² - x - w = 0, then the value of w is:
----------------------
If x-4 is a factor, 4 is a root.
----
Solve:: f(4) = 0
4^2 - 4 - w = 0
12-w = 0
w = 12
-------------
Cheers,
Stan H.
------------